Visible light has wavelengths in the range of 400 nm to 780 nm. Calculate the range of energy of the photons of visible light.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. λ base 1 = 400 nm to λ base 2 = 780 nm E = hv = hc/λ h = 6.63 * 10^-34 j – s, c = 3 * 10^8 m/s, λ base 1 = 400 nm, λ base 2 = 780 nm E base 1 = 6.63 * 10^-34 * 3 * 10^8/400 * 10^-9 = 6.63 *3/4 * 10^-19 = 5 * 10^-19 J. E base 2 = 6.63 *3/47.8 * 10^-19 = 2.55 * 10^-19 J So, the range is 5 * 10^-19 J to 2.25 * 10^-19 J.